OpenStudy (anonymous):

solve the system of equations: 5x+-y=-3 x+-4y=9

OpenStudy (anonymous):

@reginad99 1) y=5x+3x 2) Y=1/4-9/4 This solves for y.

OpenStudy (anonymous):

multiply equation 1 with -4 on both sides so we get -20x+4y=12 and add to equation 2 x-4y=9 we get -19x=21 or x=-21/19 and y=-48/19

OpenStudy (anonymous):

idk i tried!

OpenStudy (anonymous):

I would do these equations using the elimination method

OpenStudy (anonymous):

@kelliegirl33 i have no idea what method i used, i just did it.

OpenStudy (anonymous):

or substitution....x = 9 - 4y, and just sub 9 - 4y in for x in the 1st equation, and then solve for x